Gaseous methane

CH4

reacts with gaseous oxygen gas

O2

to produce gaseous carbon dioxide

CO2

and gaseous water

H2O

. What is the theoretical yield of water formed from the reaction of

1.1g

of methane and

1.9g

of oxygen gas?
Answer #1

Ans :

The balanced reaction is given as :

CH4 + 2O2 = CO2 + 2H2O

Number of mol of methane = mass / molar mass

= 1.1 g / 16.04 g/mol = 0.0686 mol

Mol oxygen gas = 1.9 g / 31.999 g/mol = 0.059 mol

1 mol methane needs 2 mol oxygen

Then 0.0686 mol methane will need : 0.0686 mol x 2 = 0.137 mol oxygen gas

So here oxygen gas is limiting reagent here

0.059 mol oxygen gas will form 0.059 mol water

So theoretical yield of water = mol x molar mass

= 0.059 mol x 18.01528 g/mol

= 1.07 g